### v1

# SynCoGen: Synthesizable 3D Molecule Generation via Joint Reaction and Coordinate Modeling

Abstract
Ensuring synthesizability in generative small molecule design remains a major challenge. While recent developments in synthesizable molecule generation have demonstrated promising results, these efforts have been largely confined to 2D molecular graph representations, limiting the ability to perform geometry-based conditional generation. In this work, we present SynCoGen (Synthesizable Co-Generation), a single framework that combines simultaneous masked graph diffusion and flow matching for synthesizable 3D molecule generation. SynCoGen samples from the joint distribution of molecular building blocks, chemical reactions, and atomic coordinates. To train the model, we curated SynSpace, a dataset containing over 600K synthesis-aware building block graphs and 3.3M conformers. SynCoGen achieves state-of-the-art performance in unconditional small molecule graph and conformer generation, and the model delivers competitive performance in zero-shot molecular linker design for protein ligand generation in drug discovery. Overall, this multimodal formulation represents a foundation for future applications enabled by non-autoregressive molecular generation, including analog expansion, lead optimization, and direct structure conditioning.
1 Introduction
Generative models significantly enhance the efficiency of chemical space exploration in drug discovery by directly sampling molecules with desired properties. However, a key bottleneck in their practical deployment is low synthetic accessibility, i.e. generated molecules are often difficult or impossible to produce in the laboratory (Gao and Coley, 2020). To address this limitation, recent work has turned to template-based methods that emulate the chemical synthesis process by constructing synthesis trees that link molecular building blocks through known reaction templates (Koziarski et al., 2024; Cretu et al., 2024; Seo et al., 2024; Gaiński et al., 2025; Gao et al., 2024; Jocys et al., 2024; Swanson et al., 2024). These representations, while useful for downstream experimental validation, do not describe the underlying 3D geometry and thus cannot capitalize on the conformational information that is often crucial for diverse chemical and biological properties.
Parallel advances in generative molecular design have explored spatial modeling at the atomic level. Inspired by advances in protein structure prediction (Yang et al., 2025; Campbell et al., 2024; Wang et al., 2025) and the development of generative frameworks such as diffusion and flow matching, recent work has focused on directly sampling 3D atomic coordinates of small molecules (Hassan et al., 2024; Jing et al., 2022; Fan et al., 2024). These methods learn to generate spatially meaningful, property-aligned conformations along with molecular graphs. The ability to model atomic coordinates directly increases the expressivity of these approaches, enabling applications such as pocket-conditioned generation (Lee and Cho, 2024), scaffold hopping (Torge et al., 2023; Yoo et al., 2024), analog discovery (Sun et al., 2025), and molecular optimization (Morehead and Cheng, 2024). However, without considering practical synthesis routes, integrating synthesizability constraints into these models remains a major challenge, and most existing 3D generative approaches do not ensure that proposed molecules can be made in practice.
This work introduces SynCoGen (Synthesizable Co-Generation), a generative modeling framework aiming to bridge the gap between 3D molecular generation and practical synthetic accessibility (Figure 1). Our main contributions are as follows:
  • Generative Framework: We propose a novel generative framework that combines masked graph diffusion with flow matching in unified time to jointly sample from the distribution over building block reaction graphs and of 3D coordinates, tying structure- and synthesis-aware modeling.
  • Molecular Dataset: We curate a new dataset SynSpace, comprising 622,766 synthesizable molecules represented as building block reaction graphs, along with 3,360,908 associated low-energy conformations. Compared to synthon-based datasets, this dataset enables models to generate more readily synthesizable molecules and directly suggest streamlined synthetic routes.
  • Empirical Validation: We demonstrate that SynCoGen achieves state-of-the-art performance in 3D molecule generation, producing physically realistic conformers while explicitly tracing reaction steps. Moreover, SynCoGen performs zero-shot 3D conditional molecular generation tasks such as linker design, highlighting its applicability for drug discovery.
Figure 1: SynCoGen is a simultaneous masked graph diffusion and flow matching model that generates synthesizable molecules in 3D coordinate space. Each node corresponds to a building block, and edges encode chemical reactions. Note that nodes are not necessarily linear and that the leaving groups are not displayed.
2 Background and Related Work
Flow Matching.
Given two distributions $\rho_{0}$ and $\rho_{1}$, and an interpolating probability path $\rho_{t}$ such that $\rho_{t=0}=\rho_{0}$ and $\rho_{t=1}=\rho_{1}$, flow matching (Lipman et al., 2023; Albergo et al., 2023; Liu et al., 2023; Peluchetti, 2023; Tong et al., 2023) aims to learn the underlying vector field $u_{t}$ that generates $\rho_{t}$. Since $u_{t}$ is not known in closed form, flow matching instead defines a conditional probability path $\rho_{t|1}$ and its corresponding vector field $u_{t|1}$. The marginal vector field $u_{t}$ can then be learnt with a parametric $v_{\theta}$ by regressing against $u_{t|1}$ with the CFM objective:
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{1}\sim\rho_{1},\mathbf{x}\sim\rho_{t|1}(\cdot|x_{1})}||{v_{t}(\mathbf{x};\theta)-u_{t|1}(\mathbf{x}|\mathbf{x}_{1})}||^{2}$ (1)
Masked Discrete Diffusion Models.
[…]
$q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x})=\begin{cases}\text{Cat}(\mathbf{z}_{s};\mathbf{z}_{t}),\hfill\mathbf{z}_{t}\neq\mathbf{m}\\ \text{Cat}(\mathbf{z}_{s};\frac{(1-\alpha_{t})\mathbf{m}+(\alpha_{s}-\alpha_{t})\mathbf{x}}{1-\alpha_{t}}),\mathbf{z}_{t}=\mathbf{m}\end{cases}$ (2)
The optimal reverse process $p_{\theta}(z_{s}\mid z_{t})$ takes the same form but with $x_{\theta}(z_{t},t)$ in place of the true $\mathbf{x}$. We adopt the zero-masking and carry-over unmasking modifications of Sahoo et al. (2024).
Multimodal Generative Models.
Multimodal data generation (e.g. text-images, audio-vision, sequences/atomic types and 3D structures) represents a challenging frontier for generative models and has seen growing interest in recent times. Current approaches for this task typically either – 1) tokenize multimodal data into discrete tokens, followed by a autoregressive generation (Meta, 2024; Xie et al., 2024; Lu et al., 2024), or 2) utilize diffusion / flow models for each modality in its native space (Lee et al., 2023; Zhang et al., 2024; Campbell et al., 2024; Irwin et al., 2025). Diffusion and flow models also offer flexibility in terms of coupled (Lee et al., 2023; Irwin et al., 2025) or decoupled (Campbell et al., 2024; Bao et al., 2023; Kim et al., 2024) diffusion schedules across modalities. SynCoGen uses a coupled diffusion schedule but at two resolutions, with discrete diffusion for graphs of building blocks and reactions, and a flow for atomic coordinates in building blocks.
3D Molecular Generation.
Several recent works (Irwin et al., 2025; Le et al., 2024; Vignac et al., 2023; Huang et al., 2023; Dunn and Koes, 2024) have studied unconditional molecular structure generation by sampling from the joint distribution over atom types and coordinates. However, these models lack the ability to constrain the design space to synthetically accessible molecules. In concurrent work, (Shen et al., 2025) uses generated 3D structures to guide GFlowNet policies in designing the graph of synthon-based linear molecules, but does not account for structural quality.
Synthesizable Molecule Generation.
Beyond directly optimizing synthesizability scores (Liu et al., 2022; Guo and Schwaller, 2025) – which are often unreliable – the predominant approach to ensuring synthetic accessibility involves modifying generative models to incorporate reaction templates. Early methods explored autoencoders (Bradshaw et al., 2019, 2020), genetic algorithms (Gao et al., 2022), and reinforcement learning (Gottipati et al., 2020; Horwood and Noutahi, 2020). More recently, GFlowNet-based (Koziarski et al., 2024; Cretu et al., 2024; Seo et al., 2024; Gaiński et al., 2025) and transformer-based (Gao et al., 2024; Jocys et al., 2024) methods have gained prominence. Such generative models have already shown practical utility in biological discovery tasks (Swanson et al., 2024). However, most methods only generate molecular graphs and do not produce 3D structures.
3 Dataset
Training a synthesizability-aware model to co-generate both 2D structures and 3D positions requires a dataset of easily synthesizable molecules in an appropriate format. In addition to atomic coordinates, this includes a graph-based representation from which plausible synthetic pathways can be inferred. A common approach is to use synthons—theoretical structural units that can be combined to form complete molecules(Baker et al., 2024; Grigg et al., 2025; Medel-Lacruz et al., 2025). Synthon-based representations do not guarantee the existence of a valid synthesis route, and they do not directly provide one even if it exists. Moreover, they lack the flexibility to constrain the reaction space, which is often critical when prioritizing high-yield, high-reliability reactions or operating within the limits of automated synthesis platforms such as self-driving labs (Abolhasani and Kumacheva, 2023).
Alternatively, many synthesis-aware generators employ external reaction simulators, such as RDKit, to couple building blocks iteratively. While convenient, such black-box steps offer no fine-grained control when a reagent has multiple reaction centers, distinct atoms or atom sets that can each serve as the specific site of bond formation or cleavage in a coupling reaction. They also do not define atom mappings between reactants and products, making it impossible to trace product atoms back to their parent building blocks, which complicates edge assignment in building block graph generation.
To overcome these limitations, we curate a new dataset SynSpace (Figure 2) comprising building block-level reaction graphs pairs with corresponding atom- and block-level graphs. We then calculate multiple 3D conformations for each graph using semi-empirical methods (Bannwarth et al., 2019). SynSpace is available at https://huggingface.co/datasets/DreiSSB/SynSpace.
Figure 2: Graphical overview of SynSpace creation process. Highly synthesizable molecules are procedurally constructed by sampling synthesis pathways from a predefined set of building blocks and reactions. Starting from an initial building block, the procedure selects a reaction center, a compatible reaction, and a suitable reactant. This process is iteratively repeated for a fixed number of reaction steps. After the final structure is assembled, multiple low-energy 3D conformations are generated.
3.1 SynSpace: Graph Generation
We begin by constructing a vocabulary of 93 commercially available, low-cost building blocks and 19 high-yield reaction templates. This vocabulary is adapted from the collection proposed by Koziarski et al. (2024), retaining reactions that (1) ensure all product atoms originate from the two input reagents, and (2) involve at most one leaving group per reagent. We emphasize that these are not only feasible chemistries, but also simple and reliable chemical reactions with readily available building blocks that can enable rapid multi-step synthesis.
We procedurally generate SynSpace from this vocabulary by iteratively coupling building block graphs at their respective reaction centers with compatible chemical reaction templates, described in detail in Section A.2. We obtained 622,766 building block reaction graphs, each constructed from 2 to 4 sequential reactions. For each resulting molecule, we generate multiple low-energy conformations and retain their atomic coordinates, with a total of 3,360,908 conformations.
Note: Injectivity.
Many commercially available building blocks contain multiple reaction centers, each compatible with a different set of corresponding reaction centers on other building blocks. In this way, a building block-level reaction graph $G_{b}=(X,E)$ is not fully specified when edges are parametrized by the reaction alone. To achieve an injective correspondence, we therefore label edges from node $i$ to $j>i$ by the triple
$e_{ij}=(r,v_{i},v_{j}),$ (3)
where $r$ is the coupling reaction and $(v_{i},v_{j})$ are the participating reaction centers on the source and destination blocks, respectively. Strictly speaking, stereoisomers that form during reactions collapse to the same $(X,E)$ representation, but this granularity suffices for the scope of the current work.
3.2 SynSpace: Conformation Generation
For each molecular graph, 50 initial conformers were generated using the ETKDG (Riniker and Landrum, 2015) algorithm (RDKit implementation). These structures were energy-minimized using the MMFF94 force field, and all conformers within 10 kcal/mol of the global minimum were retained. The resulting geometries were then re-optimized with the semi-empirical GFN2-xTB (Bannwarth et al., 2019) method, after which the same 10 kcal/mol energy threshold was applied. At every stage, redundant structures were removed by geometry-based clustering ($\text{RMSD}<1.5\AA$). This workflow yields, on average, 5.4 distinct conformers per graph. Relative to exhaustive approaches such as CREST (Pracht et al., 2024), the workflow is several orders of magnitude faster; despite occasionally omitting some conformations, the retained structures are diverse and reproduce the bond-length, bond-angle, and dihedral-angle distributions observed in CREST-derived datasets (see Section 5.1).
4 Methods
Notation.
Let $\mathcal{B}$ be the building-block vocabulary and $\mathcal{R}$ the set of reaction templates, with cardinalities $B:=|\mathcal{B}|$ and $R:=|\mathcal{R}|$. We write $N$ for the maximum number of building blocks that any molecule in the training set can contain, and $M$ for the maximum number of atoms in a single building block. For each block $b\in\mathcal{B}$ we denote its set of reaction-center atoms by $\mathcal{V}(b)$; the global maximum of these counts is $V_{\max}:=\max_{b\in\mathcal{B}}\lvert\mathcal{V}(b)\rvert$.
Hence, tensor shapes contain factors such as $B+1$ (to accommodate the masked token $\pi_{X}$ in $X$), $R\,V_{\max}^{2}+2$ (to accommodate the no-edge and masked tokens $\lambda_{E}$ and $\pi_{E}$), together with the bounds $N$ and $M$ introduced above. For any coordinate tensor $C$ and binary mask $S$ we define the mask-weighted centroid and its centered version by $\bar{C}_{S}:=\frac{\sum S\odot C}{\sum S},\qquad\tilde{C}:=C-\bar{C}_{S}.$
SynCoGen.
SynCoGen generates building block-level reaction graphs and coordinates. Each molecule is represented by a triple $(X,E,C)$ where $X\in\{0,1\}^{N\times|\mathcal{B}|+1}$ encodes the sequence of building-block identities, $E\in\{0,1\}^{N\times N\times|\mathcal{R}|V_{\max}^{2}+2}$ labels the coupling reaction (and centers) between every building block pair, and $C\in\mathbb{R}^{N\times M\times 3}$ stores all atomic coordinates. We detail the parameterization of graphs $(X,E)$ in Section B.1. Training combines two diffusion schemes: 1) a discrete absorbing process on $(X,E)$ using the categorical forward kernel of Sahoo et al. (2024), and 2) a continuous, visibility-aware process on $C$ whose endpoints are (i) a rototranslationally-aligned isotropic Gaussian and (ii) a re-centered ground truth, considering all "visible" atoms in the prior (see Section 4.2). The code is available at https://github.com/andreirekesh/SynCoGen.
4.1 Model Architecture
At each timestep $t$, SynCoGen predicts building block logits $L_{t}^{X},L_{t}^{E}$ and a shifted coordinate estimate $\hat{\tilde{C}}^{\,t}_{0}$. The total loss is the weighted sum of the cross-entropy term $\mathcal{L}_{\mathrm{graph}}$ on $(X,E)$, the masked coordinate MSE term $\mathcal{L}_{\mathrm{MSE}}$, and the short-range pairwise distance term $\mathcal{L}_{\mathrm{pair}}$ (see Section B.3 for the training algorithm and Section B.11 for loss details). We implement a modified version of SemlaFlow (Irwin et al., 2025), a $SE(3)$ equivariant architecture originally designed for all-atom molecular design, as the principal backbone to generate both coordinates and graphs. As SemlaFlow handles atom and atom-level bond representations as both inputs and outputs, we define additional building-block-to-atom featurization protocols (Section B.2) and atom-to-building-block output layers (Section B.5).
4.2 Noising Schemes
Graph Noising.
We corrupt true graphs $(X_{0},E_{0})$ using the posterior described in Section 2. In practice, as all true edge matrices $E_{0}$ are symmetric, we symmetrize the sampled probabilities for the noising and denoising of $E_{t}$ correspondingly (see Section B.6).
Coordinate Noising.
For any time $t$ where $X_{t}$ contains a masked building block, we cannot distinguish any of its $M$ possible atoms from padding. For this reason, we design a visibility mask $S_{t}$ that considers all $M$ atoms for each noisy coordinate at time $t$ as valid. We then center the prior by its visibility-masked centroid $\tilde{C_{1}}=C_{1}-\bar{C_{1}}_{S_{t}}$. Here, all atoms $a\in\operatorname{supp}(S_{t})\setminus\operatorname{supp}(S_{0})$ are potentially valid at time $t$, but represent padding indices in the true molecule.
We thus must construct and interpolate a data-prior pair $(\tilde{C}_{1},\tilde{C_{0}})$ that contains a consistent number of valid atoms $|S_{t}|$ by which both $\tilde{C}_{0}$ and $\tilde{C}_{1}$ are centered. To handle atoms that do not appear in $C_{0}$, we record their points in $\tilde{C}_{1}$, re-center $C_{0}$ by the same visibility-masked centroid, then copy the atoms to their respective indices in $C_{0}$. This process is formalized in Algorithm 1.
Here, $\mathcal{A}_{i}$ is the set of all atom indices $a$ that constitute true atoms in $X_{0}$. Essentially, we task the model with rearranging the true atoms while disregarding padding by learning to fix padding atoms in place. Note that $S_{t}=S_{0}$ for all $t$ where $X_{t}$ contains no masked building blocks.
Algorithm 1 PairData$\bigl{(}C_{0},S_{0},C_{1},t,X_{t}\bigr{)}$
Input: $C_{0}$ (clean coordinates), $S_{0}$ (atom mask), $C_{1}$ (prior sample), $t\!\in\![0,1]$, $X_{t}$ (partially masked nodes)
Output: $\tilde{C}_{0}$ (re-centered ground truth), $C_{t}$ (interpolated noisy coords)
1:$\mathcal{D}_{t}\leftarrow\{\,i\mid X_{t}[i]\neq\pi_{X}\}$triangleright denoised blocks
2:$S_{t}[i,a]\leftarrow\mathbf{1}[i\notin\mathcal{D}_{t}\lor a\in\mathcal{A}_{i}]$triangleright visibility
3:$\tilde{C}_{1}\leftarrow C_{1}-\bar{C_{1}}_{S_{t}}$
4:$\tilde{C}_{0}\leftarrow\textsc{ZeroTensor}()$
5:for all $(i,a)$ do
6:     if $S_{0}[i,a]=1$ then
7:         $\tilde{C}_{0}[i,a]\leftarrow C_{0}[i,a]-\bar{C_{1}}_{S_{t}}$
8:     else if $S_{t}[i,a]=1$ then $\triangleright$ dummy atom
9:         $\tilde{C}_{0}[i,a]\leftarrow\tilde{C}_{1}[i,a]$
10:     end if
11:end for
12:$C_{t}\leftarrow(1-t)\,\tilde{C}_{0}+t\,\tilde{C}_{1}$
13:return $\bigl{(}\tilde{C}_{0},\,C_{t}\bigr{)}$
Note: Non-Equivariance.
This process results in data pairings where both $C_{0}$ and $C_{t}$ are properly centered according to atoms that are possibly valid at time $t$. It is important to note that under this scheme, while the model is $SE(3)$-equivariant with respect to the system defined by the partial mask $S_{t}$, it is not equivariant with respect to the orientation of the molecule itself unless $\mathcal{D}_{t}^{c}\;=\;\varnothing$, as the presence and temporary validity of masked dummy atoms offsets the true atom centering and thus breaks both translational and rotational equivariance.
Flexible Atom Count.
Most 3D molecule generation methods require specifying the number of atoms during inference. Because the prior of SynCoGen is over building blocks, we naturally handle a flexible number of atoms during generation and model any excessive atoms as ghost atoms.
4.3 Training-time Constraints
SynCoGen inherits training-time simplifications from MDLM (Sahoo et al., 2024), including zero masked logit probabilities and carry-over logit unmasking. In addition, we implement the following:
  1. 1.
    No-Edge Diagonals. We set the diagonals of all edge logit predictions $L_{\theta}^{E}$ to no-edge, as no building block has a coupling reaction-induced bond to itself.
  2. 2.
    Edge Count Limit. Let
    $k_{t}:=\sum_{1\leq i<j\leq n}\mathds{1}\bigl{[}E_{t}[i,j,\cdot]\notin\{\pi_{E},\lambda_{E}\}\bigr{]},$ (4)
    be the number of unmasked true edges in the upper triangle of $E_{t}$. If $k_{t}=n-1$, we have the correct number of edges for a molecule containing $n$ building blocks and therefore set all remaining edge logits to $\lambda_{E}$.
  3. 3.
    Compatibility Masking. Assume that for some $E_{t}$ an edge entry is already denoised, $E_{t}[i,j,\cdot]=(r,v_{i},v_{j})$, meaning that building block $i$ reacts with building block $j$ via reaction $r$ and centers $v_{i}\in\mathcal{V}(X_{i})$, $v_{j}\in\mathcal{V}(X_{j})$. Define the sets of center-matched reagents
    $\textstyle\mathcal{B}_{r,v}^{A}$ :=$:=\{\,b\in\mathcal{B}\mid(b,v)\text{ matches reagent A in }r\},$ (5)
    $\mathcal{B}_{r,v}^{B}$ :=$:=\{\,b\in\mathcal{B}\mid(b,v)\text{ matches reagent B in }r\}.$
    For every node slot $i$ (resp. $j$) we construct a $|\mathcal{B}|$-dimensional binary mask
    $\textstyle\mathcal{X}_{i,k}=\mathds{1}[b_{k}\in\mathcal{B}_{r,v_{i}}^{A}],$ (6)
    $\mathcal{X}_{j,k}=\mathds{1}[b_{k}\in\mathcal{B}_{r,v_{j}}^{B}],$
    $k=1,\dots,|\mathcal{B}|.$
    so that the soft-max for $X_{t}[i,\cdot]$ (resp. $X_{t}[j,\cdot]$) is evaluated only over the 1-entries of $\mathcal{X}_{i}$ (resp. $\mathcal{X}_{j}$). Analogously, once a node identity $X_{t}[j]=b$ is denoised, incoming edge channels $(i,j)$ with $j>i$ are masked to reactions $e=(r,v_{i},v_{j})$ such that $b\in\mathcal{B}_{r,v_{i}}^{B}$.
Put simply, we restrict logits to disallow loops (e.g. macrocycles), to impose a limit on the number of edges, and to better ensure the selection of chemically compatible building blocks and reactions.
4.4 Sampling
Sampling begins by drawing a building block count $n\sim\operatorname{Cat}(\pi_{\text{frag}})$, setting the node and edge tensors to the masked tokens, $X_{1}[i,\!\cdot]=\pi_{X},\;E_{1}[i,j,\!\cdot]=\pi_{E}$ for every $0\leq i,j<N$, and padding all $(i\geq n)$ rows/columns with the no–edge token $\lambda_{E}$. The initial coordinates are an isotropic Gaussian $C_{1}\sim\mathcal{N}(0,I)^{N\times M\times 3}$. From this state, the sampler walks backwards in diffusion time, and at each step it (i) recenters the current coordinates by the visibility mask $S_{t}$ derived from $X_{t}$, (ii) generates node, edge and coordinate predictions with the trained model, (iii) applies the single-parent edge constraint (Algorithm 2), (iv) draws the next discrete state, and (v) updates coordinates by an Euler step. After a final, deterministic pass, we calculate $\bigl{(}\hat{X}_{0},\hat{E}_{0}\bigr{)}=\arg\max_{k}L_{\theta}^{E}[\cdots,k]$ and center the coordinates to yield the molecule $(\hat{X}_{0},\hat{E}_{0},\hat{C}_{0})$. Complete pseudocode is provided in Section B.4. We note our discrete and continuous schemes share a unified time. Lastly, we find inference annealing on the coordinates (see Section D.1) yields small performance gains at sampling time.
Constraint-Aware Edge Pruning.
By construction, a molecule containing $n$ connected building blocks contains exactly $n-1$ edges, and fragment $j>0$ has a unique parent $i<j$. Consequently, let $E^{t}_{\theta}\in[0,1]^{n\times n\times|\mathcal{R}|V_{\max}^{2}}$ be the soft-max edge probabilities produced at step $t$. The routine below resolves the unique parent for every building block column $j>0$ and returns a probability tensor $\tilde{E}^{t}_{\theta}$ with exactly one non–zero entry per column.
Algorithm 2 SampleEdges$\bigl{(}E^{t}_{\theta},n\bigr{)}$
Input: edge probabilities $E^{t}_{\theta}$
Output: pruned probabilities $\tilde{E}^{t}_{\theta}$
1:$\tilde{E}^{t}_{\theta}\leftarrow\mathbf{0}$
2:for $j=1$ to $n-1$ do
3:     $(i_{j},e_{j})\sim\operatorname{Cat}\!\bigl{(}\{E^{t}_{\theta}[i,j,e]\mid 0\leq i<j\}\bigr{)}$
4:     $\tilde{E}^{t}_{\theta}[i_{j},j,e_{j}]\leftarrow 1$
5:end for
6:return $\tilde{E}^{t}_{\theta}$
$\tilde{E}^{t}_{\theta}$ is then symmetrized and fed to the discrete reverse sampler described in Section B.6.
5 Experiments
5.1 De Novo 3D Molecule Generation
We first study SynCoGen in unconditional molecule generation jointly with 3D coordinates and reaction graphs. We evaluate SynCoGen against several recently published all-atom generation frameworks which produce 3D coordinates, including SemlaFlow (Irwin et al., 2025), EQGAT-Diff (Le et al., 2024), MiDi (Vignac et al., 2023), JODO (Huang et al., 2023), and FlowMol (Dunn and Koes, 2024). To isolate modeling effects from data, we retrain SemlaFlow on atomic types/coordinates in SynSpace for the same number of epochs as SynCoGen.
For each model, we sample 1000 molecules and compute stringent metrics capturing chemical soundness, synthetic accessibility, conformer quality, and distributional fidelity. Pertaining to the molecular graphs, we report the RDKit sanitization validity (Valid.) and retrosynthetic solve rate (AiZynthFinder (Genheden et al., 2020) (AiZyn.) and Syntheseus (Maziarz et al., 2025) (Synth.)). For conformers, we introduce two physics-based metrics: the median non-bonded interaction energies per atom via the forcefield method GFN-FF and via the semiempirical quantum chemistry method GFN2-xTB Bannwarth et al. (2019); Spicher and Grimme (2020); we also check PoseBusters (Buttenschoen et al., 2024) validity rate (PB). We evaluate the diversity (Div.) as the average pairwise Tanimoto similarity, novelty (Nov.) as the percentage of candidates not appearing in the training set, and the Fréchet ChemNet Distance (Preuer et al., 2018) (FCD) on the distance between generated samples and the training distribution. Details on how the metrics are computed can be found in Section D.2.
Table 1: Comparison of generative methods for de novo 3D molecule generation. ↑/↓ indicate that higher/lower is better.
[…]
Table 2: Ablation studies. We incrementally remove inference annealing, auxiliary losses, self-conditioning, scaled-noise, and constraints to see the performance difference. All results shown are at 50 epochs rather than 100 epochs in Table 1. See Sections 4.4 and 4.3, (Sections B.9, B.3 and B.11).
Method Valid. $\uparrow$ GFN-FF $\downarrow$
Base 93.5 4.871
- Inference annealing 93.5 4.933
- Auxiliary losses 85.3 5.194
- Self-conditioning 69.0 6.424
- Scaled noise 70.4 5.091
- Constraints 42.4 67.006
The results are presented in Table 1, and example molecules can be seen in Figure 8. For chemical reasonableness, we see that SynCoGen generates almost entirely valid molecules. Our generation process mimics a multi-step reaction pathway, and as a result, our molecules are significantly more synthesizable compared to baseline methods. Because AiZynthFinder and Syntheseus solve only 50–70 % of known drug-like molecules, our 50–72 % scores likely underestimate true synthesizability.
For structural reasonableness, the generated conformers reproduce the overall energy distributions and have very favorable non-covalent interaction energies as evaluated by relatively accurate computational chemistry methods, especially when compared to the baseline methods (Tables 1 and 3). This is also evident from the lack of structural changes upon further geometric relaxation (Figure 9). The low non-bonded energies indicate SynCoGen learns to sample many intramolecular interactions (Figure 8). Quantitatively, 87% of these conformers pass PoseBusters pose plausibility checks. Furthermore, SynCoGen reproduces the delicate data distribution of bond lengths, angles, and dihedrals (Figures 3 and 7). For example, SynCoGen generates fewer $sp^{2}$-$sp^{2}$ bonds that are too short, captures sharp bond angle distributions (e.g., $sp^{3}$-$sp^{3}$-$sp^{3}$), and replicates the broad dihedral angle distribution of the flexible $sp^{3}$-$sp^{3}$-$sp^{3}$-$sp^{3}$ while still covering the rigid dihedral angles (e.g., $sp^{3}$-$sp^{2}$-$sp^{2}$-$sp^{2}$). Wasserstein-1 distances and JSD can be found in Table 4, and additional plots on other bond lengths/angles/dihedrals can be found in Figure 7.
Figure 3: Conformer geometry and energy distribution comparisons. Distributions of a) bond lengths, b) bond angles, c-d) dihedral angles, e) average per-atom GFN2-xTB non-covalent interaction energies and f) average per-atom GFN-FF non-bonded interaction energies. Solid curves denote training data densities; lower subpanels in (a-d) show deviations between generated samples and data.
Training-, sampling-time and architectural ablations are presented in Tables 2 and D.1. The largest performance enhancement originates from chemistry-aware graph constraints and self-conditioning, with small contributions from other training/sampling details. The performance gap between SynCoGen and SemlaFlow retrained on SynSpace underscores that our training procedure — rather than the architecture or dataset — is the primary driver of performance.
The multi-modal model can further perform other tasks; for example, given randomly assembled molecules from reaction graphs, SynCoGen can perform zero-shot conformer generation at a quality similar to ETKDG as implemented in RDKit (Table 5).
Finally, SynCoGen captures the training distribution as indicated by the low FCD, while generally producing novel molecules. The generated samples have slightly lower diversity as a trade-off for using a (limited) set of reaction building blocks. All generated samples are unique.
5.2 Molecular Inpainting for Fragment Linking
Figure 4: Molecular inpainting. a) Fragment linking starts from three experimentally identified ligands in the PDB that contain substructure matches with our building blocks. We show three examples of linkers generated by SynCoGen per structure and the distribution of Vina docking scores. b) Proposed synthesis pathway for molecule (1) sampled from our model and c) structure of (1) (blue) docked onto PDB 7N7X using AlphaFold3 compared against the PDB ligand (beige).
To demonstrate the practical usefulness of SynCoGen, we study fragment linking (Bancet et al., 2020) to design easily synthesizable analogs of hard-to-make drugs. Fragment linking in drug design enables the construction of potent molecules, such as PROTAC bivalent protein degraders Békés et al. (2022); Chirnomas et al. (2023), by connecting smaller fragments that are known to bind distinct regions of a target site. We formulate fragment linking as a molecular inpainting task, where we fix the identity and coordinates of two fragments in a known ligand and sample its missing parts consistent with both geometry and reaction grammar.
As a case study, we pick several FDA-approved, hard-to-synthesize small molecules with experimental crystal structures, each bound to a different target protein. We select human plasma kallikrein (PDB: 7N7X), multidrug-resistant HIV protease 1 (PDB: 4EYR), and human cyclin-dependent kinase 6 (PDB: 5L2S), where each structure is complexed with a ligand that contains at least two of our building blocks. At sampling time, we condition on the substructure match by keeping fixed fragments denoised and replacing their coordinates with the corresponding linearly interpolated configuration at each time step $t$ (see Section B.13 for details).
We evaluate our generated molecules using AutoDock Vina (Figure 4(Eberhardt et al., 2021). SynCoGen consistently produces molecules with docking scores on par with or better than the native ligand while satisfying constraints on the presence of specific building blocks. AlphaFold3 (Abramson et al., 2024b) folding predictions on selected protein-ligand pairs show similar binding positions in the selected pockets as well (Figures 4 and 10). Crucially, unlike existing approaches (Schneuing et al., 2024; Igashov et al., 2024), the model links fragments using building blocks and reactions to ensure streamlined synthetic routes of the designs (Table 6). Sample multi-step synthetic pathways are shown in Figure 11.
We emphasize that using SynCoGen for fragment-linking does not require retraining. We benchmarked SynCoGen against the state-of-the-art, purpose-built fragment-linking model DiffLinker (Igashov et al., 2024). While SynCoGen produces fewer valid molecules, the valid ones have similar PoseBuster validity rate, lower intramolecular interaction energies, and no disconnected fragments. Most importantly, the retrosynthesis solve rate for DiffLinker is 0% whereas for SynCoGen, it is 58-79% (Table 6). This proof-of-concept inpainting setup is directly applicable for tasks such as scaffold hopping, synthesizable analog generation, or PROTAC design.
6 Conclusion
In this work, we introduced SynCoGen, a unified generative model that jointly samples building-block reaction graphs and atomic coordinates, thereby designing synthesizable molecules directly in Cartesian space. To enable training, we curated SynSpace, a new dataset of readily synthesizable molecules paired with low-energy 3D conformations. SynCoGen achieves state-of-the-art performance across 3D molecular generation benchmarks, while natively returning a tractable synthetic route for each structure. The practical utility of SynCoGen is demonstrated in molecular inpainting to design easily synthesizable drug analogs with geometry constraints.
With SynCoGen, several future directions in synthesizability-aware 3D generation are possible. Beyond chemical graphs and conformers, property-conditioned generation—either through multimodal guidance or direct conditioning on binding pockets—could potentially yield new, high-affinity ligands. Additionally, the current constrained set of reactions and building blocks deliberately favor reliable, automated synthesis at the expense of breadth, and its expansion would open larger regions of chemical space. Beyond drug design, the same co-generative principle may be extended to materials discovery, where synthesizability and 3D accuracy are equally pivotal.
Acknowledgments and Disclosure of Funding
The authors thank Joey Bose (University of Oxford), Stephen Lu (McGill University), and Francesca-Zhoufan Li (Caltech) for helpful discussions. This work is enabled by high-performance computing at the Digital Research Alliance of Canada and Mila.
Appendix A Chemistry Details
A.1 Building Blocks and Reactions
[…]
A.2 Graph Generation
Helper definitions.
We annotate each building block with its reaction center atom indices $\mathcal{V}(b)\subseteq V(b)$ and its and each intrinsic atom-level graph by $H(b)\;:=\;\bigl{(}V(b),\,L(b)\bigr{)}$, where $V(b)$ is the set of atoms in $b$ and $L(b)\subseteq V(b)\times V(b)$ is the set of covalent bonds internal to the block. Each reaction template $r$ is annotated with a Boolean tuple $\bigl{(}(l_{A}(r),\,l_{B}(r)\,\bigr{)}\in\{0,1\}^{2}$ describing whether reagent $A$ or reagent $B$ in $r$, respectively, contains a leaving atom.
Given the current atom graph $G_{a}=(V_{a},L_{a})$ and an atom $v\in V_{a}$ of degree 1, the routine $\textsc{UniqueNeighbor}(v)$ returns the single atom $u\in V_{a}$ such that $(u,v)\in L_{a}$. Throughout the vocabulary, every leaving-group center has exactly one neighbour.
A reaction template $r$ is considered compatible with $(b_{i},v)$ and $(\tilde{b},\tilde{v})$ if it queries for first and second reagent substructures that match $(b_{i},v)$ and $(\tilde{b},\tilde{v})$, respectively.
Figure 5: List of building blocks, their respective reaction centers (in red), and their canonical SMILES representation.
Figure 6: List of chemical reactions used to connect building blocks and their SMARTS representation. Newly formed bonds are highlighted in pink.
Algorithm 3 Fragment-by-fragment assembly with Couple
Inputs: vocab $\mathcal{B}$, reactions $\mathcal{R}$, depth limit $T$
Output: atom graph $G_{a}$, building block graph $G_{f}=(X,E)$
1:function Couple($G_{a},\;b_{i},\;\tilde{b},\;r,\;(v_{i},\tilde{v})$)
2:     append all atoms and bonds of $H(\tilde{b})$ to $G_{a}$triangleright 1. Handle leaving groups
3:     if $l_{A}(r)=1$ then $\triangleright$ $v_{i}$ leaves in reagent A
4:         $u_{i}\leftarrow$ UniqueNeighbor($v_{i}$)
5:         delete atom $v_{i}$ (and its bond) from $G_{a}$
6:         $v_{i}\leftarrow u_{i}$triangleright reroute to neighbour
7:     end if
8:     if $l_{B}(r)=1$ then $\triangleright$ $\tilde{v}$ leaves in reagent B
9:         $u_{\text{t}}\leftarrow$ UniqueNeighbor($\tilde{v}$)
10:         delete atom $\tilde{v}$ (and its bond) from $G_{a}$
11:         $\tilde{v}\leftarrow u_{\text{t}}$triangleright reroute to neighbour
12:     end if
13:     add covalent bond between $v_{i}$ and $\tilde{v}$triangleright 2. Add the cross-bond
14:     return $G_{a}$
15:end function
16:$b_{0}\leftarrow\text{UniformPick}(\mathcal{B})$; $G_{a}\leftarrow H(b_{0})$; $G_{f}\leftarrow(b_{0})$
17:for $t=1$ to $T$ do
18:     $L\leftarrow$ enumerate compatible 5-tuples $\langle b_{i},v,r,\tilde{b},\tilde{v}\rangle$
19:     if $L=\varnothing$ then break
20:     end if
21:     $(b_{i},v,r,\tilde{b},\tilde{v})\leftarrow\text{UniformPick}(L)$
22:     $e\leftarrow(r,v,\tilde{v})$
23:     $G_{a}\leftarrow\textsc{Couple}(G_{a},\,b_{i},\,\tilde{b},\,r,(v,\tilde{v}))$
[…]
25:end for
26:return $(G_{a},G_{f})$
Appendix B Method Details
B.1 Building Block-Level Representations
Let $X\in\{0,1\}^{N\times|\mathcal{B}|+1}$ be a one-hot matrix where the $i^{\text{th}}$ row encodes the identity of the $i^{\text{th}}$ building block, and let $E\;\in\;\{0,1\}^{N\times N\times|\mathcal{R}|V_{max}^{2}+2},$ where $V_{max}=\max_{b}|\mathcal{V}(b)|$. A non-zero entry $E_{ijr(v_{i},v_{j})}=1$ signals that block $i$ (center $v_{i}$) couples to block $j$ (center $v_{j}$) via reaction $r$. Graphs $(X,E)$ belonging to molecules containing $n<N$ building blocks are padded to $N$.
Reserved Channels.
We reserve a dedicated masked (absorbing) token in both vocabularies:
$\pi_{X}\in\{0,1\}^{|\mathcal{B}|},\qquad\pi_{E}\in\{0,1\}^{|\mathcal{R}|\,V_{\max}^{2}},$ (7)
where $\pi_{X}$ (resp. $\pi_{E}$) is the one-hot vector whose single 1-entry corresponds to the masked node (resp. edge) channel. Besides the masked channel, we keep a dedicated no-edge channel, encoded by the one-hot vector
$\lambda_{E}\in\{0,1\}^{|\mathcal{R}|\,V_{\max}^{2}},$ (8)
so every edge slot may take one of three mutually exclusive states: a concrete coupling label, the no-edge token $\lambda_{E}$, or the masked token $\pi_{E}$.
B.2 Atom-Level Representations
The SemlaFlow [Irwin et al., 2025] architecture propagates and updates invariant and equivariant features at the atom level. To ensure consistency with this framework, we calculate for each input graph ($X_{t},E_{t}$) atom-level one-hot atom and bond features. Crucially, these features must be flexible to arbitrary masking present in $X_{t}$ and $E_{t}$. With this in mind we set each atom feature $X^{atom}_{t}[i,a]$ to a concatenation of one-hot encodings
[…]
where $\delta_{\mathrm{sym}}(i,a)$ is the one-hot vector over possible atom types (C, N, O, B, F, Cl, Br, S, [MASK]) and ring$(i,a)$ denotes whether or not the atom is a member of a ring. Similarly, we calculate a bond feature matrix
$E^{\text{atom}}_{t}[a_{i},a_{j}]=\begin{cases}\delta_{\mathrm{order}}(a_{i},a_{j}),&\text{bond is present},\\[6.0pt] \mathbf{0}_{5},&\text{otherwise}.\end{cases}$ (10)
where $\delta_{\mathrm{order}}(a_{i},a_{j})$ is the one-hot tensor over possible bond orders (single, double, triple, aromatic, [MASK]) between $a_{i}$ and $a_{j}$. $E^{atom}_{t}$ is populated by loading the known bonds and respective bond orders within denoised building blocks. If some building block $X_{i}$ is noised, all edges between its constituent atoms $E^{atom}_{t}[i:i+M,i:i+M]$ are set to the masked one-hot index. For graphs ($X_{t},E_{t}$) corresponding to valid molecules in which all nodes and edges are denoised, we simply obtain the full bond feature matrix from the molecule described by ($X_{t},E_{t}$).
B.3 Training Algorithm
Algorithm 4 Training step for SynCoGen
$t\sim\mathcal{U}(0,1)$
2:$(X_{t},E_{t})\leftarrow q_{t}(X_{0},E_{0})$
3:$C_{1}\sim\mathcal{N}(0,I)$
4:$(\tilde{C}_{0},\tilde{C}_{t})\leftarrow\textsc{Pair}(C_{0},S_{0},C_{1},t,X_{t})$triangleright center and interpolate coordinates (Algorithm 1)
5:$(L^{X}_{t},L^{E}_{t},\hat{\tilde{C}}_{0}^{\,t})\leftarrow f_{\theta}(X_{t},E_{t},\tilde{C}_{t},n,t)$
6:$\mathcal{L}\leftarrow\mathcal{L}_{\text{graph}}+\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{pair}}$triangleright total loss (Section B.11)
7:$\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}$
B.4 Sampling Algorithm
Algorithm 5 Sampling procedure for SynCoGen
1:$n\!\sim\!\text{Cat}(\pi_{\text{frag}})$;  $(X_{1},E_{1})\!\leftarrow(\pi_{X},\pi_{E})$;  $S_{1}[i,a]\!\leftarrow\!\mathbf{1}[i<n]$triangleright draw $n$, initialize masks
2:$C_{1}\!\sim\!\mathcal{N}(0,I)$;  $\tilde{C}_{1}\!\leftarrow\!C_{1}-\bar{C}_{1,S_{1}}$triangleright center Gaussian prior by initial mask
3:for $t=1$ down to 00 do
4:     $\tilde{C}_{t}\!\leftarrow\!C_{t}-\bar{C}_{t,S_{t}}$;
5:     $(L^{X}_{t},L^{E}_{t},\hat{\tilde{C}}_{0}^{\,t})\!\leftarrow f_{\theta}(X_{t},E_{t},\tilde{C}_{t},n,t)$
6:     $\tilde{L}^{E}_{t}\!\leftarrow\!\textsc{SampleEdges}(L^{E}_{t},n)$triangleright enforce one parent per building block (Algorithm 2)
7:     $X_{t-\Delta t}\!\leftarrow\!\textsc{CatSample}(L^{X}_{t})$; $E_{t-\Delta t}\!\leftarrow\!\textsc{CatSample}(\tilde{L}^{E}_{t})$triangleright take reverse step (Section B.6)
8:     $C_{t-\Delta t}\!\leftarrow\!C_{t}+\Delta t(\hat{\tilde{C}}_{0}^{\,t}-\tilde{C}_{t})$
9:     $(X_{t},E_{t},C_{t},S_{t})\!\leftarrow(X_{t-\Delta t},E_{t-\Delta t},C_{t-\Delta t},S_{t-\Delta t})$
10:end for
11:$(L^{X},L^{E},\hat{\tilde{C}}_{0})\!\leftarrow\!f_{\theta}(X_{0},E_{0},\tilde{C}_{0},n,0)$triangleright final deterministic denoise ($t=0$)
12:$\hat{X}_{0}\!\leftarrow\!\arg\max_{k}L_{\theta}^{X}[\cdots,k];\,$ $\hat{E}_{0}\!\leftarrow\!\arg\max_{k}L_{\theta}^{E}[\cdots,k];\,$ $\hat{C}_{0}\!\leftarrow\!\hat{\tilde{C}}_{0}-\bar{\hat{\tilde{C}}}_{0,S_{0}}$
13:return $(\hat{X}_{0},\hat{E}_{0},\hat{C}_{0})$
B.5 Building Block Logit Predictions
The SemlaFlow[Irwin et al., 2025] backbone outputs atom–atom edge features $E^{\mathrm{atom}}_{\theta}\in\mathbb{R}^{B\times(NM)\times(NM)\times d_{\text{edge}}}$. To obtain building block-level tensors, we apply two parallel 2-D convolutions (one for nodes, one for edges) with stride $M$, followed by MLP classifiers that map the pooled features back to their original one-hot vocabularies.
Stride-pooled convolution.
Let $d_{edge}$ be the latent edge feature dimension. Each stream uses the block
[…]
so every $M\times M$ atom patch collapses to a single building block entry. This produces
$X_{\text{pool}}\in\mathbb{R}^{B\times d_{\text{edge}}\times N},\qquad E_{\text{pool}}\in\mathbb{R}^{B\times d_{\text{edge}}\times N\times N}.$ (12)
Node head.
We flatten $X_{\text{pool}}$ along its channel axis, concatenate the residual building block one-hot matrix $X_{t}$, and pass the result through a two-layer MLP to obtain
$L_{\theta}^{X_{t}}\in\mathbb{R}^{B\times N\times|\mathcal{B}|}.$ (13)
Edge head.
We concatenate $E_{\text{pool}}$ with the residual building block-edge one-hot tensor $E_{t}$, apply an analogous two-layer MLP, and symmetrize to produce
$L_{\theta}^{E_{t}}\in\mathbb{R}^{B\times N\times N\times|\mathcal{R}|V_{\max}^{2}}.$ (14)
Atom Features.
The SemlaFlow[Irwin et al., 2025] backbone additionally outputs atom-level node features $X^{\mathrm{atom}}_{\theta}\in\mathbb{R}^{B\times(NM)\times d_{\text{node}}}$, which are incorporated into $E^{\mathrm{atom}}_{\theta}$ via a bond refinement message-passing layer. We find that extracting both building block and edge logits directly from the refined features $E^{\mathrm{atom}}_{\theta}$ marginally improves performance relative to separately predicting $L_{\theta}^{X_{t}}$ from $X^{\mathrm{atom}}_{\theta}$ and $L_{\theta}^{E_{t}}$ from $E^{\mathrm{atom}}_{\theta}$.
B.6 Discrete Noising Scheme
Following [Sahoo et al., 2024], we adopt an absorbing (masked) state noising scheme for $X_{0}$ and $E_{0}$:
$q(X_{t}\mid X_{0})=\operatorname{Cat}\!\bigl{(}X_{t};\,\alpha_{t}X_{0}+(1-\alpha_{t})\pi_{X}\bigr{)},\qquad q(E_{t}\mid E_{0})=\operatorname{Cat}\!\bigl{(}E_{t};\,\alpha_{t}E_{0}+(1-\alpha_{t})\pi_{E}\bigr{)}.$ (15)
where $(\alpha_{t})_{t\in[0,1]}$ is the monotonically decreasing noise schedule introduced in Section 2.
Reverse categorical posterior.
For node identities, we have
$q\!\left(X_{s}\mid X_{t},\,X_{0}\right)=\begin{cases}\operatorname{Cat}(X_{s};X_{t}),&X_{t}\neq\pi_{X},\\[6.0pt] \displaystyle\operatorname{Cat}\bigl{(}X_{s};\frac{(1-\alpha_{s})\pi_{X}+\alpha_{s}\,X_{\theta}^{t}}{1-\alpha_{t}}\bigr{)},&X_{t}=\pi_{X},\end{cases}\!$ (16)
and, analogously, for edge labels
$q\!\left(E_{s}\mid E_{t},\,E_{0}\right)=\begin{cases}\operatorname{Cat}(E_{s};E_{t}),&E_{t}\neq\ \pi_{E}\\[6.0pt] \displaystyle\operatorname{Cat}\bigl{(}E_{s};\frac{(1-\alpha_{s})\pi_{E}+\alpha_{s}\,E_{\theta}^{t}}{1-\alpha_{t}}\bigr{)},&E_{t}=\pi_{E},\\[12.0pt] \end{cases}\!$ (17)
where $s<t$. Equations 16 and 17 are the direct translation of the reverse denoising process described by [Sahoo et al., 2024] into SynCoGen’s node–edge representation.
B.7 Noise Schedule Parameterization
Following MDLM [Sahoo et al., 2024], we parameterize the discrete noising schedule via $\alpha_{t}=e^{-\sigma(t)}$, where $\sigma(t):[0,1]\to\mathbb{R}^{+}$. In all experiments, we adopt the linear schedule:
$\sigma(t)=\sigma_{\max}t,$ (18)
where $\sigma_{\max}$ is a large constant; we use $\sigma_{\max}=10^{8}$ as in the original MDLM setup.
Edge Symmetrization.
After drawing the upper-triangle entries of the one-hot edge tensor $E_{s}$ in either the forward or reverse (de)noising process, we enforce symmetry by copying them to the lower triangle:
$E_{s,jie}\;=\;E_{s,ije},\qquad 0\leq i<j<n,\;\;e\in\mathcal{R}V_{\max}^{2}.$
B.8 Positional Embeddings
Though SemlaFlow[Irwin et al., 2025] is permutationally invariant by design with respect to atom positions, SynCoGen dataset molecules require that atom order be fixed and grouped by building block for reconstruction purposes. To enforce this during training, we intentionally break permutation invariance by generating and concatenating to each input coordinate sinusoidal positional embeddings representing both global atom index and building block index.
B.9 Hyperparameters
We train SynCoGen for 100 epochs with a batch size of 128 and a global batch size of 512. Note that SemlaFlow and Midi are trained for 200 epochs, and EQGAT-diff is trained for up to 800 epochs. All models are trained with a linear noise schedule (see Section B.7), with the SUBS parameterization enabled. During training, a random conformer for each molecule is selected, then centered and randomly rotated to serve as the ground-truth coordinates $C_{0}$. All atomic coordinates are normalized by a constant $Z_{c}$ describing the standard deviation across all training examples. For the pairwise distance loss $\mathcal{L}_{pair}$, we set $d$ to 3Å, adjusted for normalization. During training, for each recentered input-prior pair $(\tilde{C}_{1},\tilde{C}_{0})$ we rotationally align $C_{1}$ to $C_{0}$. When training with noise scaling and the bond loss time threshold, we set the noise scaling coefficient to 0.2 and the time threshold to 0.25, above which bond length losses are zeroed. When training with auxiliary losses, we set the weights for the pairwise, sLDDT, and bond length loss components to 0.4, 0.4, and 0.2, respectively.
B.10 Computational Resources Used
We train all models on 2 H100-80GB GPUs.
B.11 Training Losses
Here, we define several loss terms that have proved useful for stabilizing training on 3-D geometry.
By default, SynCoGen is trained with $\mathcal{L}_{\text{MSE}}$ and ${L}_{\text{pair}}$ as coordinate losses.
For a prediction $\bigl{(}L_{\theta}^{X_{t}},\,L_{\theta}^{E_{t}},\,\hat{\tilde{C}}_{0}^{\,t}\bigr{)}=f_{\theta}\!\bigl{(}X_{t},E_{t},\tilde{C}_{t},n,t\bigr{)}$, $X_{\theta}^{t}=\operatorname{softmax}\!\bigl{(}L_{\theta}^{X_{t}}\bigr{)},\;E_{\theta}^{t}=\operatorname{softmax}\!\bigl{(}L_{\theta}^{E_{t}}\bigr{)}$:
Graph loss.
Let $X_{0}$ and $E_{0}$ be the clean node and edge tensors. Following the MDLM implementation [Sahoo et al., 2024], we weigh the negative log-likelihood at step $t$ by
$w_{t}=\frac{\Delta\sigma_{t}}{\exp(\sigma_{t})-1},\qquad\Delta\sigma_{t}=\sigma_{t}-\sigma_{t-1},\;\;\sigma_{0}=0,$ (19)
where $\sigma_{t}$ is the discrete noise level. The discrete (categorical) loss is then
$\mathcal{L}_{\text{graph}}=w_{t}\bigl{(}-\!\!\log X_{\theta}^{t}[X_{0}]-\!\!\log E_{\theta}^{t}[E_{0}]\bigr{)},$ (20)
i.e. the cross-entropy between the one-hot ground truth and the predicted distributions for both nodes and edges.
MSE loss.
Let $S_{0}\in\{0,1\}^{N\times M}$ mask the atoms that exist in the clean structure and $C_{t}$ be the noisy coordinates. Denote $\mathcal{A}_{S_{0}}=\bigl{\{}(i,a):S_{0}[i,a]=1\bigr{\}}$.
$\mathcal{L}_{\text{MSE}}=\frac{1}{|\mathcal{A}_{S_{0}}|}\sum_{(i,a)\in\mathcal{A}_{S_{0}}}\bigl{\|}\hat{C}_{0}[i,a]-C_{0}[i,a]\bigr{\|}_{2}^{2},$ (21)
Pairwise loss.
$\mathcal{L}_{\text{pair}}=\sum_{\begin{subarray}{c}(i,a)<(j,b)\\ \|C_{0}[i,a]-C_{0}[j,b]\|_{2}\leq d\end{subarray}}S_{0}[i,a]\,S_{0}[j,b]\;\bigl{(}\|\hat{C}_{0}[i,a]-\hat{C}_{0}[j,b]\|_{2}-\|C_{0}[i,a]-C_{0}[j,b]\|_{2}\bigr{)}^{2},$ (22)
where $d$ is the distance cut-off for pairwise terms. The default total loss value for the model is therefore
$\mathcal{L}_{\text{{{SynCoGen}}}}=\mathcal{L}_{\text{graph}}+\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{pair}}.$ (23)
Smooth-LDDT loss [Abramson et al., 2024a].
Let $d_{ij}^{0}:=\lVert C_{0}[i]-C_{0}[j]\rVert_{2}$ and $d_{ij}^{\text{pred}}:=\lVert\hat{C}_{0}[i]-\hat{C}_{0}[j]\rVert_{2}$ be ground-truth and predicted inter-atomic distances, respectively. For each pair of atoms within a $15$ cutoff in the reference structure, we compute the per-pair score
$\operatorname{sLDDT}_{ij}=\frac{1}{4}\sum_{k=1}^{4}\sigma\!\bigl{(}\tau_{k}-\lvert d_{ij}^{\text{pred}}-d_{ij}^{0}\rvert\bigr{)},\quad\bigl{[}\tau_{1},\tau_{2},\tau_{3},\tau_{4}\bigr{]}=[0.5,1,2,4]\text{ \AA},$
where $\sigma(x)=1/(1+e^{-x})$ is the logistic function. The smooth-LDDT loss averages $1-\operatorname{sLDDT}_{ij}$ over all valid pairs,
$\mathcal{L}_{\text{sLDDT}}=\frac{\displaystyle\sum_{i<j}\mathds{1}[d_{ij}^{0}<15]\,S_{0}[i]\,S_{0}[j]\,\bigl{(}1-\operatorname{sLDDT}_{ij}\bigr{)}}{\displaystyle\sum_{i<j}\mathds{1}[d_{ij}^{0}<15]\,S_{0}[i]\,S_{0}[j]}.$ (24)
Bond-length loss.
Given a set of intra-fragment bonds $\mathrm{bonds}=\{(p,q)\}$ extracted from the vocabulary, we penalize deviations in predicted bond lengths:
$\mathcal{L}_{\text{bond}}=\frac{1}{|\mathrm{bonds}|}\sum_{(p,q)\in\mathrm{bonds}}\bigl{|}\lVert\hat{C}_{0}[p]-\hat{C}_{0}[q]\rVert_{2}-\lVert C_{0}[p]-C_{0}[q]\rVert_{2}\bigr{|}.$ (25)
Self-Conditioning.
The modified SemlaFlow [Irwin et al., 2025] backbone operates on node and edges features at the atomic level, but outputs unnormalized prediction logits $\hat{X_{0}}\in\{0,1\}^{N\times|\mathcal{B}|}$ and $\hat{E_{0}}\in\;\{0,1\}^{N\times N\times|\mathcal{R}|V_{max}^{2}}$. We therefore implement modified self-conditioning for SynCoGen that projects previous step graph predictions $\hat{X_{0}}_{cond}$ and $\hat{E_{0}}_{cond}$ to the shape of $X^{atom}_{t}$ and $E^{atom}_{t}$ using an MLP.
B.12 Conformer generation
We randomly assembled 50 molecules with the reaction graph and used the standard conformational search (iMTD-GC) in CREST with GFN-FF to find all reference conformers. For both SynCoGen and RDKit ETKDG, we sampled 50 conformers per molecule and computed the coverage and matching scores. We used a relatively strict RMSD threshold of $\tau=0.75\ \mathrm{\AA}$.
Formally, COV is defined as:
$\mathrm{COV}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\!\biggl{[}\min_{1\leq j\leq M}\mathrm{RMSD}(m_{i},\,g_{j})\leq\tau\biggr{]},$ (26)
where $\mathbf{1}[\cdot]$ is the indicator function, $m_{i}$ are the $N$ generated conformers and $g_{j}$ are the $M$ reference conformers. And MAT is defined as:
$\mathrm{MAT}=\frac{1}{N}\sum_{i=1}^{N}\min_{1\leq j\leq M}\mathrm{RMSD}(m_{i},\,g_{j}).$ (27)
B.13 Molecular inpainting
For the inpainting experiments in Section 5.2, we keep two fragments $\mathcal{D}=\{\mathcal{D}^{(1)},\mathcal{D}^{(2)}\}$ and their coordinates fixed and sample the remaining part of the molecule. We follow Section B.4 and initialize the graph prior $X_{1}$ with the one-hot encoding of the desired fragment $i$ at a specified node index in the graph (decided at random or based on the structure of the original molecule, so that it matches its scaffold). For each denoised fragment $\mathcal{D}^{(i)}$, we replace its coordinates at each time $t>0.03$ during sampling by
$C_{t}^{(i)}\;=\;(1-t)\,\tilde{C}_{0}^{(i)}\;+\;t\,\tilde{C}_{1}^{(i)},$
where $\tilde{C}_{0}^{(i)}$ and $\tilde{C}_{1}^{(i)}$ are the centered ground-truth and prior coordinates of fragment $i$, respectively, and all other fragments are updated as shown in Section B.4. For any $t\leq 0.03$, which for 100 sampling steps amounts to the last three steps in the path, we follow normal Euler steps as shown in Section B.4 to allow a refinement of the fixed coordinates in line with the rest of the predicted ones for the rest of the fragments. We empirically observed that this led to molecules with lower average energies.
Appendix C Baseline comparisons.
For all baselines, we sampled 1000 molecules with random seeds on an A100 GPU and reported averaged results over three runs.
SemlaFlow
We evaluated SemlaFlow using the sampling script in the official codebase on GitHub [^1] https://github.com/rssrwn/semla-flow/, available under the MIT License. We reported results for a model trained on the GEOM [Axelrod and Gomez-Bombarelli, 2022] dataset (by sampling from the checkpoints provided in the repository) and from a model trained on our dataset (see Table 1). We trained SemlaFlow using the default hyperparameters for 150 epochs on a single conformer per molecule.
EQGAT-diff, MiDi, JODO, FlowMol
We evaluated EQGAT-diff, Midi, JODO, using their official implementations provided on GitHub [^2] https://github.com/jule-c/eqgat_diff/, https://github.com/cvignac/MiDi, https://github.com/GRAPH-0/JODO, https://github.com/Dunni3/FlowMol, available under the MIT License. We modified the example sampling script to save molecules as outputted from the reverse sampling, without any post-processing. For MiDi, we evaluated the uniform model. For FlowMol, both CTMC and Gaussian models were evaluated and reported.
Appendix D Extended results and discussion
D.1 Sampling Ablations
By default, SynCoGen implements a linear noise schedule and samples for 100 timesteps. To evaluate the effect of step count and noise schedule choice on performance, we provide experiments with step count decreased to 50 and 20, as well as modified noising to follow a log-linear and geometric schedule. All results listed subsequently can be assumed to use the default noise schedule and step count.
We additionally follow FoldFlow to implement inference annealing, a time-dependent scaling on Euler step size that was found to empirically improve in-silico results in protein design Bose et al. [2024]. We studied multiplying the Euler step size at time $t$ by $5t$, $10t$, and $50t$. In practice, we employ $10t$ for our experiments unless otherwise noted.
We find that noising and de-noising building blocks according to a linear noise schedule generally achieves good performance, which during inference sees most unmasking occur in the final steps. An aggressive denoising schedule for the discrete fragments yields significantly worse validity (Geometric and Loglinear). Inference annealing that speeds up continuous denoising in the beginning but slows it down near the end helps to inform discrete unmasking and can slightly improve discrete generation validity, energies, and PoseBusters validity. As a sanity check to evaluate whether simultaneous generation is necessary for good performance using SynCoGen, we evaluate an inference configurations where all building blocks and reactions are noised until a single final prediction step (FinalOnly) where we find performance using the default parameters to be superior.
Table 3: Sampling ablations. Results are averaged over 1000 generated samples, except retrosynthesis solve rate (out of 100). All results shown are at 50 epochs rather than 100 epochs in Table 1.
[…]
D.2 Metrics
We here describe metric computation details that are absent in the main text.
For synthesizability evaluation, we used the public AiZynthFinder and Syntheseus models. Due to the speed of these models, we only evaluate 100 randomly sampled generated examples. For AiZynthFinder, we used the USPTO policy, the Zinc stock, and we extended the search time to 800 seconds with an iteration limit of 200 seconds. For Syntheseus, we used the LocalRetro model with Retro* search under default settings, with Enamine REAL strict fragments as the stock. We additionally appended our building blocks as the stock, but found no meaningful difference in solved rates, presumably as most of our building blocks are already in the utilized stock. We note that we replaced all boranes with boronic acids due to simplifications made in our modeling (see Section A.2).
For energy evaluation, all results are from single-point calculations. For GFN-FF, we report the total energy minus the bond energies (equivalent to the sum of angle, dihedral, bond repulsion, electrostatic, dispersion, hydrogen bond, and halogen bond energies) as the intramolecular non-bond energies, and average it over the number of atoms. For GFN2-xTB, we report the dispersion interaction energies as the intramolecular non-covalent energies. We note that the total energies and bonded energies follow very similar trends. We note that MMFF94 energies are not parameterized for boron; therefore, we report them only for the Wasserstein distances in Table 4 and inpainting task in Table 6. Figures 3 and 7 show distributions obtained from 1,000 molecules generated by each generative method, along with 50,000 subsampled molecules from their respective training datasets. Gaussian kernel density estimation (bandwidth = 0.15) was used for linear distributions, while von Mises kernel density estimation ($\kappa=25$) was applied for circular distributions. Wasserstein-1 distances (computed linearly for lengths and energies, and on the circle for angles and dihedrals) were calculated using the Python Optimal Transport Package [Flamary et al., 2021].
D.3 De novo 3D molecule generation
Additional sample molecules and their synthetic pathways can be found in Figures 8 and 11. We further take a subset of random molecules and perform GFN2-xTB geometry optimization, and find that our generated samples closely recapitulate the optimized geometries (Figure 9). We also report the Wasserstein-1 distances and Jensen-Shannon divergence for the generative models on bond lengths, angles, and energies (Table 4), and show additional statistics in Figure 7. Lastly, we show the model’s ability to generate conformers with given graphs instead of co-generation in  Table 5.
Figure 7: Additional conformer bond length, angle, and dihedral distribution comparisons. a-b) Bond lengths, c-d) bond angles, e-f) dihedral angles. Solid curves denote training data densities; lower subpanels in (a-d) show deviations between generated samples and data.
Figure 8: Unconditionally sampled random molecules from  SynCoGen.
Figure 9: A subset of randomly sampled molecules from  SynCoGen and further optimized by GFN2-xTB until convergence. Alignment RMSD is shown below the molecular structures.
Table 4: Wasserstein-1 distance ($W_{1}$) and Jensen–Shannon divergence (JSD) for the generative models (lower is better). For bond lengths, angles, and dihedrals, we computed the average $W_{1}$ and JSD for the top 10 prevalent lengths/angles/dihedrals. Comparisons are made to the respective training set.
(a) Bond dihedrals
Method $W_{1}$ JSD
SynCoGen $7.01$ $0.29$
SemlaFlow SynSpace $6.50$ $0.22$
SemlaFlow $7.76$ $0.28$
EQGAT-Diff $8.48$ $0.29$
MiDi $9.32$ $0.38$
JODO $5.47$ $0.31$
FlowMol-CTMC $13.69$ $0.35$
FlowMol-Gauss $18.85$ $0.46$
(b) Bond angles
Method $W_{1}$ JSD
SynCoGen $1.36$ $0.22$
SemlaFlow SynSpace $1.64$ $0.28$
SemlaFlow $1.18$ $0.21$
EQGAT-Diff $1.37$ $0.16$
MiDi $1.41$ $0.21$
JODO $0.59$ $0.12$
FlowMol-CTMC $1.90$ $0.24$
FlowMol-Gauss $3.68$ $0.30$
(c) Bond lengths
Method $W_{1}$ JSD
SynCoGen $0.0171$ $0.34$
SemlaFlow SynSpace $0.0320$ $0.48$
SemlaFlow $0.0200$ $0.38$
EQGAT-Diff $0.0039$ $0.13$
MiDi $0.0142$ $0.31$
JODO $0.0034$ $0.12$
FlowMol-CTMC $0.0089$ $0.20$
FlowMol-Gauss $0.0152$ $0.28$
(d) GFN2-xTB non-covalent $E$
Method $W_{1}$ JSD
SynCoGen $0.0838$ $0.33$
SemlaFlow SynSpace $0.0125$ $0.16$
SemlaFlow $0.0249$ $0.16$
EQGAT-Diff $0.0073$ $0.12$
MiDi $0.0084$ $0.14$
JODO $0.0031$ $0.11$
FlowMol-CTMC $0.0605$ $0.26$
FlowMol-Gauss $0.0322$ $0.19$
(e) GFN-FF non-bonded $E$
Method $W_{1}$ JSD
SynCoGen $1.37$ $0.28$
SemlaFlow SynSpace $1.09$ $0.22$
SemlaFlow $1.52$ $0.16$
EQGAT-Diff $1.69$ $0.18$
MiDi $1.80$ $0.19$
JODO $1.33$ $0.12$
FlowMol-CTMC $1.53$ $0.17$
FlowMol-Gauss $2.13$ $0.17$
(f) MMFF total $E$
Method $W_{1}$ JSD
SynCoGen $6.59$ $0.089$
SemlaFlow SynSpace $54.63$ $0.22$
SemlaFlow $69.56$ $0.24$
EQGAT-Diff $4.80$ $0.076$
MiDi $19.00$ $0.11$
JODO $22.07$ $0.11$
FlowMol-CTMC $41.95$ $0.15$
FlowMol-Gauss $26.96$ $0.14$
Table 5: With given reaction graphs, comparison of mean coverage (COV) and matching accuracy (MAT) for RDKit ETKDG and zero-shot conformer generation using SynCoGen.
Method COV (%) $\uparrow$ MAT (Å) $\downarrow$
RDKit 0.692 0.657
SynCoGen 0.614 0.693
D.4 Molecular inpainting experiments
Three protein–ligand complexes (PDB IDs 7N7X [^3] https://www.rcsb.org/structure/7N7X, 5L2S [^4] https://www.rcsb.org/structure/5L2S and 4EYR [^5] https://www.rcsb.org/structure/4EYR) were selected for molecular inpainting of the ligand structures. These ligands were chosen because they are prominent FDA-approved drugs, and they are typically challenging to synthesize, but the key functional groups are present in our building blocks. Specifically, 4EYR contains ritonavir, a prominent HIV protease inhibitor on the World Health Organization’s List of Essential Medicines; 5L2S contains abemaciclib, an anti-cancer kinase inhibitor that is amongst the largest selling small molecule drugs; 7N7X contains berotralstat, a recently approved drug that prevents hereditary angioedema. Note that for 4EYR, the inpainting was done using the ligand geometry from the PDB entry 3NDX [^6] https://www.rcsb.org/structure/3NDX, but docking was performed with 4EYR because the protein structure in 3NDX contained issues – nonetheless, both entries contain the same protease and ligand.
In addition to the experiments in Section 5.2, we evaluate SynCoGen’s conditional sampling performance for the fragment linking framework against the state-of-the-art model DiffLinker [Igashov et al., 2024]. While DiffLinker is trained for fragment-linking, our model performs zero-shot fragment linking without any finetuning. For both models, the size of the linker was chosen so that it matches that of the original ligand: 2 extra nodes were sampled for SynCoGen and 15 linking atoms for DiffLinker in the case of 5L2S, while 3 extra nodes and 25 linking atoms were sampled for 4EYR and 7N7X. We specified leaving groups (for SynCoGen) and anchor points (for DiffLinker) so that the fragments are linked at the same positions as in the ligand. Results are shown in Table 6. No retrosynthetic pathways were found for the molecules in DiffLinkers, while SynCoGen models synthetic pathways and synthetic pathways can be easily drawn, with examples for 4EYR shown in Figure 11. This out-of-distribution task for SynCoGen leads to fewer valid molecules; however, for the valid candidates, SynCoGen has lower interaction energies and achieves 100% connectivity as it uses reaction-based assembly, whereas DiffLinker can sample disconnected fragments.
Table 6: Molecular inpainting task. Results are averaged over 1000 generated samples, except retrosynthesis solve rate (out of 100).
Method Target AiZyn. $\uparrow$ Synth. $\uparrow$ Valid. $\uparrow$ Connect. $\uparrow$ MMFF $\downarrow$ GFN-FF $\downarrow$ GFN2-xTB $\downarrow$ Diversity $\uparrow$ PB $\uparrow$
DiffLinker 5L2S 0 0 95.8 95.09 14.22 7.52 -0.95 0.60 49.3
4EYR 0 0 93.7 81.86 20.01 8.49 -1.03 0.81 35.0
7N7X 0 0 95.8 74.65 20.51 7.99 -1.09 0.78 37.5
SynCoGen 5L2S 73 79 57.6 100 10.11 6.77 -0.78 0.62 27.3
4EYR 72 58 46.9 100 12.80 6.58 -0.86 0.64 32.0
7N7X 53 69 50.6 100 4.243 6.60 -0.80 0.67 56.1
(a) 5L2S
(b) 3NDX/4EYR
(c) 7N7X
Figure 10: Structural overlays of the native protein (gray) and its native ligand (blue) with AlphaFold3-predicted folds of a subset of generated ligands (green) for (a) 5L2S, (b) 3NDX/4EYR, and (c) 7N7X.
Figure 11: Synthetic pathways for molecules generated in the molecular inpainting task for target 3NDX/4EYR. The final product is shown in blue, and the inpainted fragments are shown in red.